# Beyond Input Activations: Identifying Influential Latents by Gradient Sparse Autoencoders

### v1

@@ Abstract
Sparse Autoencoders (SAEs) have recently emerged as powerful tools for interpreting and steering the internal representations of large language models (LLMs). However, conventional approaches to analyzing SAEs typically rely solely on input-side activations, without considering the causal influence between each latent feature and the model’s output. This work is built on two key hypotheses: (1) activated latents do not contribute equally to the construction of the model’s output, and (2) only latents with high causal influence are effective for model steering. To validate these hypotheses, we propose Gradient Sparse Autoencoder (GradSAE), a simple yet effective method that identifies the most influential latents by incorporating output-side gradient information. Our code is available at https://anonymous.4open.science/r/sae_gradient-10FF.
@@ 2.1 Problem Statement
Let $\mathcal{V}$ be the vocabulary, $\mathcal{X}\in\mathcal{V}^{N}$ an input sequence, and $\mathcal{Y}\in\mathcal{V}^{M}$ the corresponding LLM-generated output. The hidden representation at layer $l$ is $\mathbf{Z}^{(l)}\in\mathbb{R}^{N\times D}$, where $D$ is the hidden dimension. We omit the superscript (l) for simplicity in the rest of the paper. A pre-trained SAE is inserted at layer $l$ with parameters $\mathbf{W}_{\text{enc}}\in\mathbb{R}^{D\times C}$ and $\mathbf{W}_{\text{dec}}\in\mathbb{R}^{C\times D}$, where $C$ is the latent dimension of SAE and $C\gg D$. Given the representation $\mathbf{Z}$ of the entry $\mathcal{X}$, SAE first decomposes $\mathbf{Z}$ as a sparse latent activation $\mathbf{H}\in\mathbb{R}^{N\times C}$, and then loselessly restores $\hat{\mathbf{Z}}\in\mathbb{R}^{N\times D}$ with $\mathbf{H}$ as:
@@ 2.2 Proposed GradSAE Framework
We illustrate our proposed GradSAE framework for estimating the influences of sparse latent activations $c=1,...,C$ in Figure 1a. Following previous work Feng et al. (2018); Wu et al. (2024), we define the influence of a certain latent $c$ at the $n$-th input token on the output $\mathcal{Y}$ as the change in prediction with and without sparse latent activation $\mathbf{H}_{n,c}\in\mathbb{R}$:
@@ 3.1.2 Perturbation Experiment (RQ1 & RQ2)
$\vspace{-0.4cm}\mathcal{Z}_{\text{high}}$ $=\text{argmax}_{\mathcal{Z}^{\prime}\subset\mathcal{Z}_{\text{NZ}},|\mathcal{Z}^{\prime}|=K}\sum_{c\in\mathcal{Z}^{\prime}}\mathbf{g}_{c},$ (4)
$\mathcal{Z}_{\text{low}}$ $=\text{argmin}_{\mathcal{Z}^{\prime}\subset\mathcal{Z}_{\text{NZ}},|\mathcal{Z}^{\prime}|=K}\sum_{c\in\mathcal{Z}^{\prime}}\mathbf{g}_{c},\vspace{-0.4cm}$ (5)

### v2

@@ Abstract
Sparse Autoencoders (SAEs) have recently emerged as powerful tools for interpreting and steering the internal representations of large language models (LLMs). However, conventional approaches to analyzing SAEs typically rely solely on input-side activations, without considering the causal influence between each latent feature and the model’s output. This work is built on two key hypotheses: (1) activated latents do not contribute equally to the construction of the model’s output, and (2) only latents with high causal influence are effective for model steering. To validate these hypotheses, we propose Gradient Sparse Autoencoder (GradSAE), a simple yet effective method that identifies the most influential latents by incorporating output-side gradient information. Our code is available at https://github.com/Tizzzzy/sae_gradient.
@@ 2.1 Problem Statement
Let $\mathcal{V}$ be the vocabulary, $\mathcal{X}\in\mathcal{V}^{N}$ an input sequence, and $\mathcal{Y}\in\mathcal{V}^{M}$ the corresponding LLM-generated output. The hidden representation at layer $l$ is $\mathbf{Z}^{(l)}\in\mathbb{R}^{N\times D}$, where $D$ is the hidden dimension. We omit the superscript (l) for simplicity in the rest of the paper. A pre-trained SAE is inserted at layer $l$ with parameters $\mathbf{W}_{\text{enc}}\in\mathbb{R}^{D\times C}$ and $\mathbf{W}_{\text{dec}}\in\mathbb{R}^{C\times D}$, where $C$ is the latent dimension of SAE and $C\gg D$. Given the representation $\mathbf{Z}$ of the entry $\mathcal{X}$, SAE first decomposes $\mathbf{Z}$ as a sparse latent activation $\mathbf{H}\in\mathbb{R}^{N\times C}$, and then restores $\hat{\mathbf{Z}}\in\mathbb{R}^{N\times D}$ with $\mathbf{H}$ as:
@@ 2.2 Proposed GradSAE Framework
We illustrate our proposed GradSAE framework for estimating the influences of sparse latent activations c=1,,Cc=1,...,C in Figure 1a. Following previous work Feng et al. (2018); Wu et al. (2024), we define the influence of a certain latent $c$ at the $n$-th input token on the output $\mathcal{Y}$ as the change in prediction with and without sparse latent activation $\mathbf{H}_{n,c}\in\mathbb{R}$:
@@ 3.1.2 Perturbation Experiment (RQ1 & RQ2)
$\vskip-11.38092pt\mathcal{Z}_{\text{high}}$ $=\text{argmax}_{\mathcal{Z}^{\prime}\subset\mathcal{Z}_{\text{NZ}},|\mathcal{Z}^{\prime}|=K}\sum_{c\in\mathcal{Z}^{\prime}}\mathbf{g}_{c},$ (4)
$\mathcal{Z}_{\text{low}}$ $=\text{argmin}_{\mathcal{Z}^{\prime}\subset\mathcal{Z}_{\text{NZ}},|\mathcal{Z}^{\prime}|=K}\sum_{c\in\mathcal{Z}^{\prime}}\mathbf{g}_{c},\vskip-11.38092pt$ (5)
@@ 5 Conclusions
Acknowledgments
Ninghao Liu is supported by the National Science Foundation (NSF) Grant #2223768 and #2507128. Mengnan Du is supported by National Science Foundation (NSF) Grant #2310261. The views and conclusions in this paper are those of the authors and should not be interpreted as representing any funding agencies.